# Prospects of phase-adaptive cooling of levitated magnetic particles in a hollow-core photonic-crystal fibre

### v2

@@ Abstract
We analyze the feasibility of cooling of classical motion of a micro- to nano-sized magnetic particle, levitated inside a hollow-core photonic crystal fiber. The cooling action is implemented by means of controlling the phase of one of the counter-propagating fiber guided waves. Direct imaging of the particle’s position, followed by the subsequent updating of the control laser’s phase leads to Stokes type of cooling force. We provide estimates of cooling efficiency and final achievable temperature, taking into account thermal and detection noise sources. Our results bring forward an important step towards using trapped micro-magnets in sensing, testing the fundamental physics and preparing the quantum states of magnetization.
@@ I Introduction
Although free space optical tweezers possess tremendous sensing capabilities, their performance is limited by diffraction and issues in the optical manipulation of trapped particles. These limitations can be addressed by trapping micro-and nanoparticles in hollow-core photonic crystal fibre (HC-PCF). This configuration offers unique advantages, including diffraction free trapping and possibility to tune the environment surrounding the particle by introducing gas, vacuum or liquid inside the hollow-core [9, 10]. Another key advantage of optical trapping in HC-PCF is the ability to guide trapped particles over distances several orders of magnitude larger than the Rayleigh length of the trapping light [11, 12, 13, 14]. This is achieved by means of an optical conveyor belt formed from the standing wave optical trap created by two counter-propagating beams. The particles can then be transported or even halted at a particular position in the fiber by changing the relative phase [15] between counter-propagating modes.
@@
Figure 1: A magnetic micro-sphere is trapped in the potential minimum of a standing wave inside a hollow-core optical photonic crystal fibre. The standing wave minima can be shifted in the $y$-direction of the fibre by the modification of the phase $\phi(t)$ of the right-propagating guided control laser. Detection of the particle’s position $y_{\text{det}}(t)$ gives an estimate $y_{\text{det}}$ of the true value $y(t)$, with an additional noise term $\eta W(t)$ modeled as a Wiener process $W(t)$ multiplied by the efficiency of the imaging scheme (poorer resolution translates into a larger $\eta$). Phase adaptive feedback results from the monitoring of the detected quadrature $y_{\text{det}}$ and the subsequent adjustment of $\phi$.
@@
We propose here an alternative route, where external optical damping is switched on during the pressure reduction stage and switched off in the next stage. This ensures the stability of the system and allows for the reach of very low pressures, where the particle is isolated from the thermal environment. The feedback loop involves a phase-adaptive mechanism [32], providing an effective Stokes-type damping force [38]. The feedback force comes from the continuous adjustment of the relative phase $\phi(t)$ between the two counter-propagating trapping lasers (see Fig. 1). This consequently allows for a controllable shifting of the potential minimum and thus the particle’s equilibrium position. In a first step, one detects light scattered to the side of the fibre and feeds the information into a motion controller. The deduced quantity is the particle’s dimensionless position $q_{\text{det}}(t)=q(t)+\eta W^{\text{det}}(t)$, which contains the exact value of the position quadrature $q(t)$, to which one adds inherent detection noise of amplitude $\eta$ and modelled as a Wiener process $W^{\text{det}}(t)$ [39]. The detected signal is then used to estimate $\phi(t)=(\mathcal{C}/\Omega)[q_{\text{det}}(t)-q_{\text{det}}(t-\tau)]/\tau$, where $\mathcal{C}$ and $\tau$ are adjustable parameters. Subsequently, the phase deduced is used to control the right propagating laser, leading to the desired optical damping effect. Notably, Grass et. al. [40] have demonstrated the trapping and cooling of a dielectric silica particle inside HC-PCF. The feedback cooling action in [40] is achieved by using an additional laser and modulating the power of the optical trap. However, phase adaptation in our technique is fundamentally different and requires only the adjustment of the relative phase, unlike power, between counter-propagating beams. One of the main advantages of phase actuation is the absence of extraneous heating that may arise from the direct intensity modulation. Another advantage of phase actuation is the controlled transportation of the cold particle from one place to another. This is not possible by simply changing the power [40].
@@ II Model and equations
Let us assume that a magnetic particle of mass $m$ is trapped in the standing wave formed by two counter-propagating waves with frequency $\omega_{\ell}$ and wave-vector ±$\pm k_{\ell}$ along the $y$-direction. The plus component of the electric field amplitude can be expressed as a superposition of $\mathcal{E}_{0}e^{ik_{\ell}y}e^{-i\omega_{\ell}t}$ and $\mathcal{E}_{0}e^{-ik_{\ell}y}e^{-i\omega_{\ell}t}e^{-2i\phi(t)}$ counterpropagating waves, where $\mathcal{E}_{0}$ and $\phi(t)$ are the amplitude and the relative phase of the light fields. In a frame rotating at $\omega_{\ell}$, the total electric field is then $\mathcal{E}(y,t)=2\mathcal{E}_{0}\cos[k_{\ell}y+\phi(t)]$. This leads to a position and time dependent potential energy $-\alpha^{\prime}\mathcal{E}(y,t)^{2}/4$ for the particle, where $\alpha^{\prime}$ represents the real part of the particle’s polarisability. This allows in turn for the computation of the gradient optical force, in the $y$-direction, to be obtained as
@@
In doing so we have reduced the description to dimensionless quadratures $y=y_{\text{zpm}}q$ and $p=p_{\text{zpm}}p$ where the zero-point motion for the position is $y_{\text{zpm}}=\sqrt{\hbar/(m\Omega)}$ (the extent of the wavefunction at zero temperature) and for the momentum $p_{\text{zpm}}=\sqrt{\hbar m\Omega}$. Here, $\hbar$ is the reduced Planck’s constant. Also, we introduced the notation $\beta=\kappa_{\text{trap}}/({k_{\ell}}p_{\mathrm{zpm}})$, with dimensions of frequency. The Wiener increment for the thermal environment has zero average and only surviving correlations $\braket{dW^{\text{th}}(t)dW^{\text{th}}(t)}=dt$ at equal times (all correlations at different times vanish). The thermal occupancy corresponding to a bath of temperature $T_{\mathrm{th}}$ (assumed large) is written as $n_{\mathrm{th}}\approx k_{B}$/${}_{\mathrm{th}}/(\hbar\Omega)$, where $k_{B}$ is the Boltzmann constant. For example, assuming $\Omega/2\pi=1~{}$kHz and $T_{\mathrm{th}}=300~{}$K, one obtains $n_{\mathrm{th}}=10^{10}$, very far from the quantum regime. From Eqs. (3), it can be easily verified that in the absence of additional actuation force, the system achieves equilibrium at a rate $\gamma$ with its thermal environment at temperature $T_{\mathrm{th}}$ (see Appendix F for details).
@@ IV Discussions
Let us address here the feasibility of such a scheme to cool the classical thermal motion of a magnetic micro-sphere levitated inside the optical hollow-core fibre. To this end we connect to the experimental setup described in Ref.  [33], making use of some of the parameters relevant for current experiments. We optimize the resulting expression for the final occupancy by taking into account the influence of gas pressure and we describe a procedure for feedback implementation. The validity of the analytical estimates is checked against numerical simulations of the stochastic difference equations.
@@ IV.1 Optimizing the final occupancy
In a setup described in Appendix C, the magnetic microparticle can be stably trapped until around $1$ mbar. However, after this it escapes from the trap due to an increase in internal bulk temperature of the particle and insufficient damping at low pressure [37]. This not only restricts the experiment to be performed at a pressure below $1$ mbar but also limits its quality factor. A possible solution to these limitations can be provided by using the phase-adaptive feedback operation in order to keep the particle stable in the optical trap during the vacuum chamber evacuation procedure. As shown in Eq. (17), the final phonon occupancy in the system is limited by the product $\eta\beta\mathcal{C}$. Experimentally, the parameter $\beta$ can be controlled by the trap stiffness which in turn can be adjusted by parameters such as the trapping power, and core size of the fibre. For the setup in Ref. [33], $\beta=\kappa_{\mathrm{trap}}/(k_{l}p_{\mathrm{zpm}})=(\sqrt{m}\Omega^{3/2}\lambda_{l})/(2\pi\sqrt{\hbar})=1.3\times 10^{11}$ Hz. The noise floor for our current detection scheme is $17$ pm/$\sqrt{\mathrm{Hz}}$ [see Appendix C] and from this we can obtain the parameter $\eta$. To do so, we take into account of the spectral contribution due to detection noise (see appendix D for details) and around $\omega\approx\Omega$, we find $\eta=9.2\times 10^{6}~{}\sqrt{\mathrm{Hz}}$. The amplifier gain $\mathcal{C}$ can be set to an optimum value in order to avoid subsequent additive noise while maintaining optimum signal-noise ratio during the stages of detection and feedback. To obtain this factor, we take $\tau\rightarrow 0$ in Eq. (4). This leads to $\phi=\mathcal{C}p_{det}\approx\mathcal{C}\sqrt{n_{\text{th}}}$, where in a thermal state the variance of the momentum quadrature is taken to be $n_{\text{th}}$. The upper bound of the detected phase is $2\pi$. Using this, we can estimate the maximum value of the gain parameter as $\mathcal{C}_{max}=2\pi/\sqrt{n_{\text{th}}}$. On the other hand, the optimum value of $\mathcal{C}$ can be obtained by minimizing the phonon occupation in Eq. (19). For $\tau\rightarrow 0$, this gives following optimum value of the amplifier gain
@@ IV.2 Numerical simulations
Figure 4: Phase adaptive feedback cooling. (a) Temporal evolution of the phonon occupancy (log-log scale) for different pressure levels. The occupancy is obtained from a numerical simulation of Eqs. (3), incorporating both thermal and detection noises, along with the pressure dependent damping behavior governed by Eq.  (21). The dotted lines show the validity of the analytically predicted final occupancy given by Eq. (17). In the absence of feedback cooling (solid red line), the system quickly thermalizes with the environment at 1 mbar, leading to high phonon-occupancy. However, with the feedback activation, the phonon occupancy is reduced, as depicted for pressures of 1 mbar (solid magenta line), 10 $\mu$bar (solid cyan line), and 0.1 $\mu$bar (solid blue line). (b) Phase-space representation of an initial thermal state (red points) and the cooling state (blue points) obtained for $p=0.1~{}\mu$bar, and (c) for a corresponding pressure, enlarged view of the thermal distribution of the final cold state. Parameters used are $\Omega/2\pi=34.5$ kHz, $T=2\pi/\Omega$, $\tau/T=7.9\times 10^{-4}$, $T_{\mathrm{th}}=300$ K, $n_{\mathrm{th}}=1.8\times 10^{8}$, $\beta/2\pi=20.6$ GHz, $\mathcal{C}=\mathcal{C}_{opt}$, $\eta=9.2\times 10^{6}\sqrt{\mathrm{Hz}}$, $\mu_{0}=18.1$ Pa.s, density of the trapped particle = 5.11 g/$\mathrm{cm}^{3}$, and $R=0.66~{}\mu$m.
@@ V Conclusions
We presented the prospects of cooling the thermal excess fluctuations linked to the mechanical motion of a magnetic microparticle, optically levitated inside a hollow-core photonic crystal fibre and in the presence of a thermal gas. The cooling is achieved through an engineered phase-adaptive feedback loop that provides an effective Stokes type viscous damping force to the system. This is accomplished by directly imaging the particle’s position followed by the adjustment of relative phase between two counter-propagating fibre guided waves. We provided an analytical estimate of the final phonon occupation taking into account of thermal as well as the measurement noises. Our numerical simulations validate the analytical results. Our findings are consequential for using trapped magnetic particles as sensors for feeble stray magnetic fields [36]. Future work, analysing the feasibility of reaching the quantum ground state, will aim at extending this procedure to prepare the non-classical states of magnetization [48, 49, 50] and testing aspect of fundamental quantum physics [51, 52].

### v3

# Prospects of phase-adaptive cooling of levitated magnetic particles in a hollow-core photonic-crystal fiber

@@ Abstract
We analyze the feasibility of cooling of classical motion of a micro- to nano-sized magnetic particle, levitated inside a hollow-core photonic crystal fiber. The cooling action is implemented by means of controlling the relative phase between counter-propagating fiber guided waves. Direct imaging of the particle’s position allows dynamic phase adjustments that produce a Stokes-type cooling force. We provide estimates of cooling efficiency and final achievable temperature, taking into account thermal and detection noise sources. Our results bring forward an important step towards using trapped micro-magnets in sensing, testing the fundamental physics and preparing the quantum states of magnetization.
@@ I Introduction
Although free space optical tweezers possess tremendous sensing capabilities, their performance is limited by diffraction and issues in the optical manipulation of trapped particles. These limitations can be addressed by trapping micro-and nanoparticles in hollow-core photonic crystal fiber (HC-PCF). This configuration offers unique advantages, including diffraction free trapping and possibility to tune the environment surrounding the particle by introducing gas, vacuum or liquid inside the hollow-core [9, 10]. Another key advantage of optical trapping in HC-PCF is the ability to guide trapped particles over distances several orders of magnitude larger than the Rayleigh length of the trapping light [11, 12, 13, 14]. This is achieved by means of an optical conveyor belt formed from the standing wave optical trap created by two counter-propagating beams. The particles can then be transported or even halted at a particular position in the fiber by changing the relative phase [15] between counter-propagating modes.
@@
Figure 1: A magnetic micro-sphere is trapped in the potential minimum of a standing wave inside a hollow-core optical photonic crystal fiber. The standing wave minima can be shifted in the $y$-direction of the fiber by the modification of the phase $\phi(t)$ of the right-propagating guided control laser. Detection of the particle’s position $y_{\text{det}}(t)$ gives an estimate $y_{\text{det}}$ of the true value $y(t)$, with an additional noise term $\eta W(t)$ modeled as a Wiener process $W(t)$ multiplied by the efficiency of the imaging scheme (poorer resolution translates into a larger $\eta$). Phase adaptive feedback results from the monitoring of the detected quadrature $y_{\text{det}}$ and the subsequent adjustment of $\phi$.
@@
We propose here an alternative route, where external optical damping is switched on during the pressure reduction stage and switched off in the next stage. This ensures the stability of the system and allows for the reach of very low pressures, where the particle is isolated from the thermal environment. The feedback loop involves a phase-adaptive mechanism [32], providing an effective Stokes-type damping force [38]. The feedback force comes from the continuous adjustment of the relative phase $\phi(t)$ between the two counter-propagating trapping lasers (see Fig. 1). This consequently allows for a controllable shifting of the potential minimum and thus the particle’s equilibrium position. In a first step, one detects light scattered to the side of the fiber and feeds the information into a motion controller. The deduced quantity is the particle’s dimensionless position $q_{\text{det}}(t)=q(t)+\eta W^{\text{det}}(t)$, which contains the exact value of the position quadrature $q(t)$, to which one adds inherent detection noise of amplitude $\eta$ and modelled as a Wiener process $W^{\text{det}}(t)$ [39]. The detected signal is then used to estimate $\phi(t)=(\mathcal{C}/\Omega)[q_{\text{det}}(t)-q_{\text{det}}(t-\tau)]/\tau$, where $\mathcal{C}$ and $\tau$ are adjustable parameters. Subsequently, the phase deduced is used to control the right propagating laser, leading to the desired optical damping effect. Notably, Grass et. al. [40] have demonstrated the trapping and cooling of a dielectric silica particle inside HC-PCF. The feedback cooling action in [40] is achieved by using an additional laser and modulating the power of the optical trap. However, phase adaptation in our technique is fundamentally different and requires only the adjustment of the relative phase, unlike power, between counter-propagating beams. One of the main advantages of phase actuation is the absence of extraneous heating that may arise from the direct intensity modulation. Another advantage of phase actuation is the controlled transportation of the cold particle from one place to another. This is not possible by simply changing the power [40].
@@ II Model and equations
Let us assume that a magnetic particle of mass $m$ is trapped in the standing wave formed by two counter-propagating waves with frequency $\omega_{\ell}$ and wave-vector ±$\pm k_{\ell}$ along the $y$-direction. The plus component of the electric field amplitude can be expressed as a superposition of $\mathcal{E}_{0}e^{ik_{\ell}y}e^{-i\omega_{\ell}t}$ and $\mathcal{E}_{0}e^{-ik_{\ell}y}e^{-i\omega_{\ell}t}e^{-2i\phi(t)}$ counterpropagating waves, where $\mathcal{E}_{0}$ and $\phi(t)$ are the amplitude and the relative phase of the light fields. In a frame rotating at $\omega_{\ell}$, the total electric field is then $\mathcal{E}(y,t)=2\mathcal{E}_{0}\cos[k_{\ell}y+\phi(t)]$. This leads to a position and time dependent potential energy $-\alpha^{\prime}\mathcal{E}(y,t)^{2}/4$ for the particle, where $\alpha^{\prime}$ represents the real part of the particle’s polarisability. For small displacement from the center of the standing wave trap, our theoretical model assumes a dipole-like gradient force on the particle’s center of mass (see Appendix C.1 for details). This allows in turn for the computation of the gradient optical force, in the $y$-direction, to be obtained as
@@
In doing so we have reduced the description to dimensionless quadratures $y=y_{\text{zpm}}q$ and $p=p_{\text{zpm}}p$ where the zero-point motion for the position is $y_{\text{zpm}}=\sqrt{\hbar/(m\Omega)}$ (the extent of the wavefunction at zero temperature) and for the momentum $p_{\text{zpm}}=\sqrt{\hbar m\Omega}$. Here, $\hbar$ is the reduced Planck’s constant. Also, we introduced the notation $\beta=\kappa_{\text{trap}}/({k_{\ell}}p_{\mathrm{zpm}})$, with dimensions of frequency. The Wiener increment for the thermal environment has zero average and only surviving correlations $\braket{dW^{\text{th}}(t)dW^{\text{th}}(t)}=dt$ at equal times (all correlations at different times vanish). The thermal occupancy corresponding to a bath of temperature $T_{\mathrm{th}}$ (assumed large) is written as $n_{\mathrm{th}}\approx k_{B}$/${}_{\mathrm{th}}/(\hbar\Omega)$, where $k_{B}$ is the Boltzmann constant. For example, assuming $\Omega/2\pi=34.5~{}$kHz and $T_{\mathrm{th}}=300~{}$K, one obtains $n_{\mathrm{th}}\sim 10^{8}$, very far from the quantum regime. From Eqs. (3), it can be easily verified that in the absence of additional actuation force, the system achieves equilibrium at a rate $\gamma$ with its thermal environment at temperature $T_{\mathrm{th}}$ (see Appendix F for details).
@@ IV Discussions
Let us address here the feasibility of such a scheme to cool the classical thermal motion of a magnetic micro-sphere levitated inside the optical hollow-core fiber. To this end we connect to the experimental setup described in Ref.  [33], making use of some of the parameters relevant for current experiments. We optimize the resulting expression for the final occupancy by taking into account the influence of gas pressure and we describe a procedure for feedback implementation. The validity of the analytical estimates is checked against numerical simulations of the stochastic difference equations.
@@ IV.1 Optimizing the final occupancy
In a setup described in Appendix C, the magnetic microparticle can be stably trapped until around $1$ mbar. However, after this it escapes from the trap due to an increase in internal bulk temperature of the particle and insufficient damping at low pressure [37]. This not only restricts the experiment to be performed at a pressure below $1$ mbar but also limits its quality factor. A possible solution to these limitations can be provided by using the phase-adaptive feedback operation in order to keep the particle stable in the optical trap during the vacuum chamber evacuation procedure. As shown in Eq. (17), the final phonon occupancy in the system is limited by the product $\eta\beta\mathcal{C}$. Experimentally, the parameter $\beta$ can be controlled by the trap stiffness which in turn can be adjusted by parameters such as the trapping power, and core size of the fiber. For the setup in Ref. [33], $\beta=\kappa_{\mathrm{trap}}/(k_{l}p_{\mathrm{zpm}})=(\sqrt{m}\Omega^{3/2}\lambda_{l})/(2\pi\sqrt{\hbar})=1.3\times 10^{11}$ Hz. The noise floor for our current detection scheme is $17$ pm/$\sqrt{\mathrm{Hz}}$ [see Appendix C] and from this we can obtain the parameter $\eta$. To do so, we take into account of the spectral contribution due to detection noise (see appendix D for details) and around $\omega\approx\Omega$, we find $\eta=9.2\times 10^{6}~{}\sqrt{\mathrm{Hz}}$. The amplifier gain $\mathcal{C}$ can be set to an optimum value in order to avoid subsequent additive noise while maintaining optimum signal-noise ratio during the stages of detection and feedback. To obtain this factor, we take $\tau\rightarrow 0$ in Eq. (4). This leads to $\phi=\mathcal{C}p_{det}\approx\mathcal{C}\sqrt{n_{\text{th}}}$, where in a thermal state the variance of the momentum quadrature is taken to be $n_{\text{th}}$. The upper bound of the detected phase is $2\pi$. Using this, we can estimate the maximum value of the gain parameter as $\mathcal{C}_{max}=2\pi/\sqrt{n_{\text{th}}}$. On the other hand, the optimum value of $\mathcal{C}$ can be obtained by minimizing the phonon occupation in Eq. (19). For $\tau\rightarrow 0$, this gives following optimum value of the amplifier gain
@@ IV.2 Numerical simulations
Figure 4: Phase adaptive feedback cooling. (a) Temporal evolution of the phonon occupancy (log-log scale) for different pressure levels. The occupancy is obtained from a numerical simulation of Eqs. (3), incorporating both thermal and detection noises, along with the pressure dependent damping behavior governed by Eq.  (21). The dotted lines show the validity of the analytically predicted final occupancy given by Eq. (17). In the absence of feedback cooling (solid red line), the system quickly thermalizes with the environment at 1 mbar, leading to high phonon-occupancy. However, with the feedback activation, the phonon occupancy is reduced, as depicted for pressures of 1 mbar (solid magenta line), 10 $\mu$bar (solid cyan line), and 0.1 $\mu$bar (solid blue line). (b) Phase-space representation of an initial thermal state (red points) and the cooling state (blue points) obtained for $p=0.1~{}\mu$bar, and (c) for a corresponding pressure, enlarged view of the thermal distribution of the final cold state. Parameters used are $\Omega/2\pi=34.5$ kHz, $T=2\pi/\Omega$, $\tau/T=7.9\times 10^{-4}$, $T_{\mathrm{th}}=300$ K, $n_{\mathrm{th}}=1.8\times 10^{8}$, $\beta/2\pi=20.6$ GHz, $\mathcal{C}=\mathcal{C}_{opt}$, $\eta=9.2\times 10^{6}\sqrt{\mathrm{Hz}}$, $\mu_{0}=18.1$ Pa.s, density of the trapped particle = 5.11 g/$\mathrm{cm}^{3}$, and $R=0.66~{}\mu$m (see Appendix C.1 for discussion on Mie scattering).
@@
In the preceding analysis, we illustrated axial (one-dimensional) feedback cooling via phase control to mitigate the dominant instability observed in experiments [33]. While the fiber mode’s intensity profile offers intrinsic transverse confinement, full trap stability requires active transverse stabilization. Parametric methods like transverse phase modulation [32] or intensity modulation [28] could extend stabilization to all three dimensions. A synergistic approach combining cold damping via phase control with phase-adaptive parametric method [32] shows potential for robust axial as well as transverse stabilization by harnessing the complementary strengths of both techniques. Furthermore, the incorporation of electrical cooling techniques [44, 45] in the radial direction may further improve system performance and robustness.
IV.3 Internal temperature of the levitated particle
@@
IV.4 Laser frequency noise effects
Let us now consider the effect of fluctuations in the trapping laser frequency, which arise due to uncertainty in the laser phase. We model the noisy laser field as having a fixed amplitude and a time-dependent fluctuating phase [50], so that $\mathcal{E}_{0}=|\mathcal{E}_{0}|e^{i\phi_{l}(t)}$, where $\phi_{l}$ is a random Gaussian variable of zero mean and variance $\sigma_{\phi_{l}}$. The effect of this phase noise (see Appendix G for details) can be captured by averaging the total electric field, $\langle\mathcal{E}(y,t)\rangle$. Using the property of a Gaussian noise, we have $\langle e^{i\phi_{l}(t)}\rangle=e^{-\sigma_{\phi_{l}}/2}$, where averaging is taken over many realizations. By applying error propagation techniques to the governing equations, we find that the phase uncertainty rescales the trapping frequency from $\Omega$ to $\Omega e^{-\sigma_{\phi_{l}}/2}$ and reduces the effective cooling rate from $\bar{\Gamma}$ to $\bar{\Gamma}e^{-\sigma_{\phi_{l}}}$. In this approach, the detection noise amplitude $\eta$ and optimized gain factor $\mathcal{C}_{opt}$ are considered independent from the phase noise contributions. The impact of laser phase noise is shown in Fig. 6 for different values of $\sigma_{\phi_{l}}$. The ensemble-averaged trajectories align precisely with the analytical predictions, demonstrating a systematic decrease in cooling rate as phase noise variance grows. This relationship is directly reflected in the negative slope of the curves shown in Fig. 6. Beyond its role in reducing the cooling rate, phase noise has only a marginal impact—even when combined with thermal and detection noise—resulting in an almost negligible change in the final phonon occupancy.
Figure 6: Effect of trapping laser phase noise on system dynamics for a pressure of $0.1~{}\mu$bar. Solid lines represent simulated trajectories for phase noise variances $\sigma_{\phi_{l}}=0$ (red), 0.3 (blue) and 0.6 (green), corresponding to trap frequency changes of $0\%$, $\sim 14\%$ and $\sim 26\%$, respectively. Dotted lines denote the analytically predicted steady-state average occupancy values. All other parameters are same as in Fig. 4.
@@ V Conclusions
We presented the prospects of cooling the thermal excess fluctuations linked to the mechanical motion of a magnetic microparticle, optically levitated inside a hollow-core photonic crystal fiber and in the presence of a thermal gas. The cooling is achieved through an engineered phase-adaptive feedback loop that provides an effective Stokes type viscous damping force to the system. This is accomplished by directly imaging the particle’s position followed by the adjustment of relative phase between two counter-propagating fiber guided waves. We provided an analytical estimate of the final phonon occupation taking into account of thermal as well as the measurement noises. Our numerical simulations validate the analytical results. Our findings are consequential for using trapped magnetic particles as sensors for feeble stray magnetic fields [36]. Future work, analysing the feasibility of reaching the quantum ground state, will aim at extending this procedure to prepare the non-classical states of magnetization [51, 52, 53] and testing aspect of fundamental quantum physics [54, 55].